A health clinic uses a solution of bleach to sterilize petri dishes in which cultures are grown. The sterilization tank contains 120 gal of a solution of 4% ordinary household bleach mixed with pure distilled water. New research indicates that the concentration of bleach should be 5% for complete sterilization. How much of the solution should be drained and replaced with bleach to increase the bleach content to the recommended level?
Answer by richwmiller(17219)

0.04*120-0.04x+1x=0.05(120)
4.8+0.96x=6
0.96x=6-4.8
0.96x=1.2
x=1.25 gal
check
0.04*118.75+1*1.25=0.05(120)
4.75+1.25=6
6.0=6
ok